Find the solution of the given initial value problem in explicit form.
y'=(1-13x)y^2 , y(0)=-1/4

I'm in differential equations and it has unfortunately been some time since I took Calc II.
it appears that I'm getting stuck near when one would integrate y'*y
Answer #1

Given y= (1 - 13 2) y, (0) = - -7 dy = (1-13a) y => 3 dy = (1-13x) dx Hene, variables aqy are separable. B . I.o. B. s, we g